OpenStudy (anonymous):

physics question, suppose two people pull on rope. person A is 50 kg and pulls towards the right (+) with 30 N force. person b who is 40 kg pulls towards the left with a 20 N force. assume the rope is massless. I dont get why the tension in the rope is the same magnitude. For the rope the forces on it are F1 - F2 = 0*a. but F1 and F2 should be the 30N and 20N , so it should be not zero but 10 N to the right for tension. something doesnt make sense here

OpenStudy (anonymous):

You are suggesting that at one part of the rope, there is 10 newtons of tension, and at the other end there is 0 tension (i.e. slack)?

OpenStudy (anonymous):

I'm not sure why you're looking at the forces on the rope. You're assuming it's massless, so you can't really talk about the rope's acceleration.

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

well i was using my book. it says the forces on the rope are F1 - F2 = 0. isnt tension the force on the rope?

OpenStudy (anonymous):

What are you being asked to find?

OpenStudy (anonymous):

A rope is one element. It can't have variable force along it. So whatever the tension is in the rope, it has to be the same all along the rope.

OpenStudy (anonymous):

Here is the full paragraph. "Suppose two people pull the ends of a rope with oppositely directed forces F1 and F2 (in bold). The rope also exerts forces -F1 and -F2 on the people (in bold)From the second law, we know F1 - F2 = ma , where m is the mass of the rope and a is its acceleration. If that acceleration is zero , or if the mass of the rope is so small that we can idealize the rope as having zero mass, then F1 - F2 = 0. In this special situation the forces exerted by the rope on the two people are equal in magntitude

OpenStudy (anonymous):

I agree with that.

OpenStudy (anonymous):

well im wondering, what is the force on the rope , if it says F1 - F2 = 0, then F1 = F2, but 30 N does not equal 20 N

OpenStudy (anonymous):

What it's actually saying is that it's impossible for person A to be pulling with 30N of force and person B to be pulling with 20N of force and for there to be zero acceleration. But you know that already since: $$F=ma$$ If F were not 0, you'd have an acceleration.

OpenStudy (anonymous):

in my example, one guy pulls 30 N right and the other guy pulls 20 N left

OpenStudy (anonymous):

I don't see where you got 30N and 20N

OpenStudy (anonymous):

i made up an example

OpenStudy (anonymous):

In your example, there would be acceleration towards the guy pulling with 30N.

OpenStudy (anonymous):

you mean 10 N

OpenStudy (anonymous):

There would be a net force of 10N which would result in an acceleration in the direction of the guy pulling with 30N force.

OpenStudy (anonymous):

Yes, but you are contradicting the hypothisis

OpenStudy (anonymous):

my question is, what is the force on the rope? and why does it say the force exerted by the rope on the two people is equal in magnitude. and which force is the tension. so there are 3 questions

OpenStudy (anonymous):

They are saying that If the rope is not accelerating, we know the forces on both ends are equal

OpenStudy (anonymous):

how do you get the forces at both ends are equal, and what is this force? the tension?

OpenStudy (anonymous):

and which force is it, what is exerting on what

OpenStudy (anonymous):

you said the forces at both ends are equal. what do you mean by that. the force that the rope exerts on the two people are equal?

OpenStudy (anonymous):

No, the force on the rope.

OpenStudy (anonymous):

but there are different forces, one is 30 N and one is 20 N

OpenStudy (anonymous):

which we stipulated

OpenStudy (anonymous):

Yes, you made up that example, where the conclusion is false, and trying to make a claim about the implication.

OpenStudy (anonymous):

the conclusion is false?

OpenStudy (anonymous):

you said the force on the rope at both ends is equal. clearly this is false, it is 20 N and 30 N , they are not equal

OpenStudy (anonymous):

So the acceleration is not 0

OpenStudy (anonymous):

wait, youre saying the net force on both ends is equal in magnitude?

OpenStudy (anonymous):

but my rope is massless

OpenStudy (anonymous):

I'm saying IF the acceleration is 0, then the NET force on the rope is 0.

OpenStudy (anonymous):

why does the book say F1-F2 = 0, so F1 = F2,

OpenStudy (anonymous):

You can't talk about force acting on massless things.

OpenStudy (anonymous):

what is F1 and F2 ?

OpenStudy (anonymous):

It says F1-F2 is 0 BECAUSE it claimed already that the acceleration is 0

OpenStudy (anonymous):

so then F1 and F2 are zero?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

is tension the F1 and F2

OpenStudy (anonymous):

\[\sum F = ma\]

OpenStudy (anonymous):

If a =0 then \[\sum F = F1-F2 = 0 \implies F1 = F2\]

OpenStudy (anonymous):

But if the forces do not sum to 0 than the acceleration of the rope will not be 0.

OpenStudy (anonymous):

yes but in my example, what is F1 and F2

OpenStudy (anonymous):

is the acceleration equal to 0?

OpenStudy (anonymous):

nomo, if you cant type , you should restart.

OpenStudy (anonymous):

www.openstudy.com start over

OpenStudy (anonymous):

A review of massless rope mechanics: A rope carries a tension. This tension is the force felt at the endpoints of the rope by whatever the rope is attached to. So, if you pull against a rope caught tied to a wall: |------ -> 30N the tension in the rope is 30N. But, by action/reaction, the wall is also pulling on the rope. And since there is no acceleration, it is pulling with an equal and opposite force of -30N. Now let's look at your example. You now have two people, pulling on the rope. One is pulling harder than the other. Without doing the math, it's obvious that whoever is pulling harder is going to move the other person. That's exactly what happens: 20N <--------> 30N --> everything moves this way So your next question is, what's the tension in the rope? Well, the tension in the rope has to be 20N, because it's the force the weaker puller is pulling with. The rest of the force of the stronger puller, 10N, goes into acceleration. So: 1. The tension in the rope is 20N. 2. The net external force on the rope is 10N. Hope this helps.

OpenStudy (anonymous):

lets see

OpenStudy (anonymous):

i get a different answer

OpenStudy (anonymous):

30 N - T = 50kg * a -20 N + T = 40 kg * a

OpenStudy (anonymous):

Heh. Cantorset, you are in fact correct.

OpenStudy (anonymous):

Those are the equations of motion, and I should have written them out before posting.

OpenStudy (anonymous):

so 10 N = 90 kg *a , so a = 1/9 m*s^-2

OpenStudy (anonymous):

Yes, I agree.

OpenStudy (anonymous):

So what is T?

OpenStudy (anonymous):

one sec, my browser is crashing

OpenStudy (anonymous):

weird, one sec

OpenStudy (anonymous):

plug back in

OpenStudy (anonymous):

so 30 - 50/9 N = T

OpenStudy (anonymous):

24.444 N is the tension

OpenStudy (anonymous):

so natho was wrong

OpenStudy (anonymous):

The problem with this exercise is that in the absence of any frictional forces, you 're going to have that both people accelerate toward each other which will remove the tension.

OpenStudy (anonymous):

so what is F1 and F2, i still dont see

OpenStudy (anonymous):

So your acceleration is not constant.

OpenStudy (anonymous):

F1 and F2 are the tension forces?

OpenStudy (anonymous):

F1 is the force of one person pulling, F2 is the force of the other.

OpenStudy (anonymous):

but thats false, because 30 - 20 is not 0

OpenStudy (anonymous):

Who said it was = 0?

OpenStudy (anonymous):

is the a = 0?

OpenStudy (anonymous):

The equations of motion that cantorset wrote are valid at the instant when both people start pushing, as are the results. However, polpak is correctly pointing out that this is a dynamic problem, not a static problem. Odds are you'll never actually find a problem like this in a problem set, because there is too much information missing to be able to give a complete answer.

OpenStudy (anonymous):

nomo, so do you agree that 24.444 is the tension?

OpenStudy (anonymous):

pol, the book answer

OpenStudy (anonymous):

In the instant when they both pull, (and in the absence of any friction) The tension is 24.444

OpenStudy (anonymous):

Is the acceleration given to be 0?

OpenStudy (anonymous):

"Suppose two people pull the ends of a rope with oppositely directed forces F1 and F2 (in bold). The rope also exerts forces -F1 and -F2 on the people (in bold)From the second law, we know F1 - F2 = ma , where m is the mass of the rope and a is its acceleration. If that acceleration is zero , or if the mass of the rope is so small that we can idealize the rope as having zero mass, then F1 - F2 = 0. In this special situation the forces exerted by the rope on the two people are equal in magnitude, and the rope can be thought of as simply transmitting a force from one person to the otherl the force at any point in the rope is referred to as tension. It is the same everywhere in the rope only if the rope is unaccelerated or if is idealized as massless

OpenStudy (anonymous):

No, but the rope is given to be massless

OpenStudy (anonymous):

So the acceleration is 0.

OpenStudy (anonymous):

But that's not a problem, it's a lesson.

OpenStudy (anonymous):

You CANNOT have forces acting on massless things.

OpenStudy (anonymous):

Your example is a problem, and it's a problem not covered by this lesson, since it clearly stipulates that the acceleration is 0; in your case, it isn't.

OpenStudy (anonymous):

hmmm, so the book is contradicting itself?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

You are trying to contradict the book

OpenStudy (anonymous):

But not making a reasoned argument.

OpenStudy (anonymous):

No, you're contradicting the book. You are constructing a situation where the acceleration cannot be zero, and trying to analyze the problem as if it could.

OpenStudy (anonymous):

as if it were 0.

OpenStudy (anonymous):

ok , i know the acceleration is not zero

OpenStudy (anonymous):

So then the book doesn't apply anymore.

OpenStudy (anonymous):

Then the forces cannot be the same.

OpenStudy (anonymous):

ok lets look at the problem again. the forces on the rope is F1 - F2 = ma , the mass of the rope is neglible. so we have F1 - F2 = 0

OpenStudy (anonymous):

Therefore the F1 = F2

OpenStudy (anonymous):

If you contradict by saying \[F1 \ne F2\] then \[ma \ne 0.\]

OpenStudy (anonymous):

the end.

OpenStudy (anonymous):

but didnt we stipulate earlier that F1 = 30 N and F2 = 20 N

OpenStudy (anonymous):

You cannot say that ma = 0 if f1 not equal f2

OpenStudy (anonymous):

The sum of the forces IS ma.

OpenStudy (anonymous):

ok, hmmmm